A 60N motor winds a cable around a drum. (I) If the cable lifts a 3-ton load of bricks to a height of 12cm in 5 sec, calculate the efficiency of the motor. (II) At what rate is work done against friction?

1
Expert's answer
2022-06-14T08:41:17-0400

Given:

"F=60\:\rm kN"

"m=3000\:\rm kg"

"h=0.12\:\rm m"

"t=5\:\rm s"


(a)

"\eta=\frac{mgh}{Fh}=\frac{mg}{F}=\frac{3000*9.8}{60000}=0.49=49\%"

(b)

"P=\frac{mgh}{t}=\frac{3000*9.8*0.12}{5}=706\rm\: W"